Sample Article for ISCA Format

I. M. Author  U. R. Coauthor
My Department  Coauthor Department
My Institute  Coauthor Institute
City, St Zip  City, ST Zip

Abstract

Please type your abstract here. There are no specific size limitations. Please be brief yet concise. It must fit within the column boundaries specified with light blue lines.

1 Introduction

Please type your introduction text here. The heading text is 12 point bold font.

1.1 Subsections

There is 1 blank line before the section heading and one afterwards. Heading text is 11 point bold font. You may cite references of the form [1] or [2], andLa-Tex will keep track of numbering. The numbers are based on the order you place them in the bibliography, not the order they appear in the text. They should be in alphabetical order. LaTex will put square brackets around the number in the text of your paper. You must run the latex process twice to allow the references to format properly. You will see a message about a missing .aux file. Simply latex it a second time and it will be fine.

2 Summary

This guide will get you through a basic article with no figures and no equations. Refer to your Latex Manual if you need to include these items.

Acknowledgements

This will show you how to do an unnumbered subsection.

References

[1] I. M. Author, An Article I Wrote,” The Journal Name, Vol. 11, pp. 3-14, 1982.

[2] Next Author, Title of Book, the Publisher, 1982.